OpenStudy (anonymous):

Stella has five times a many books as Tina, if Stella gave Tina 16 books, they would have the same amount. How many books did each girl have???

OpenStudy (geometry_hater):

8 and 40

OpenStudy (geometry_hater):

8 * 5 = 40 Check 40 - 16 = 24 8 +16 = 24

OpenStudy (anonymous):

why did you multiply 8*5?

OpenStudy (geometry_hater):

because 8 is what tina started with the problem says that stella has 5 times as much so we multiply 8 times 5 to get 40 thats how many books stella has

OpenStudy (anonymous):

where did you get 8 from?

OpenStudy (geometry_hater):

well actually i kinda have to guess for it a little you see if it was 9 it wouldnt have worked

OpenStudy (geometry_hater):

err if we guessed 9 it wouldnt have worked

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Solve the following for stella and tina: Yes, one does not always have to stick with x and y for unknowns. {stella = 5 tina, stella - 16 = tina + 16} {stella = 40, tina = 8}

OpenStudy (anonymous):

so how would you solve algebraically?